two numbers equal 92 the difference between the two numbers is 21
Answer by stanbon(75887)

two numbers equal 92 the difference between the two numbers is 21
---------------
EQUATIONS:
x+y = 92
x-y = 21
------------------
Add the two equations.
2x = 113
x = 56.5
y = 92-56.5 = 35.5
=====================
Cheers,
Stan H.